OpenStudy (anonymous):

A car travels 300 miles in 5 hours. Assuming the speed is constant, how long does it take the car to travel 82 miles?

OpenStudy (saifoo.khan):

use the short method, Ratio Method.

OpenStudy (saifoo.khan):

Speed : Hours 300 : 5 82 : x Cross multiply.

OpenStudy (saifoo.khan):

300x = 82 * 5 \[x = \frac{410}{300}\] x = Answer.

OpenStudy (anonymous):

thanks

OpenStudy (saifoo.khan):

welks

OpenStudy (anonymous):

1 hour and 22 minutes

OpenStudy (anonymous):

could you show your workings?

OpenStudy (anonymous):

sorry, 300 divided by 5 = 60, 82-60=22, 60 + 22=82, 82 minutes is 1 hour and 22 minutes